# CharacterFactory: Sampling Consistent Characters with GANs for Diffusion Models

### v1

@@ 1. Introduction
In the evolving realm of text-to-image generation, diffusion models have emerged as indispensable tools for content creation (rombach2022high, ; chen2023pixart, ; zhang2023text, ). However, the inherent stochastic nature of the generation models leads to the inability to generate consistent subjects in different contexts directly, as shown in Figure 1. Such consistency can derive many applications: illustrating books and stories, creating brand ambassador, movie making, developing presentations, art design, identity-consistent data construction and more.
@@
Subject-driven methods work by either representing a user-specific image as a new word (li2023photomaker, ; wang2024stableidentity, ; chen2023dreamidentity, ) or learning image feature injection (wang2024instantid, ; ye2023ip, ; wei2023elite, ) for consistent image generation. Their training paradigms typically include per-subject optimization and encoder pretraining on large-scale datasets. The former usually requires lengthy optimization for each subject and tends to overfit the appearance in the input image (gal2022image, ; ruiz2023dreambooth, ). The latter consumes significant computational costs and struggles in stably capturing the identity and its details (li2023photomaker, ; wang2024instantid, ). However, these methods attempt to produce images with the same identity as the reference images, instead of creating a new character in various contexts. A feasible way is that a text-to-image model is used in advance to create a new character’s image and then subject-driven methods are adopted to produce images with consistent identity. Such a two-stage workflow could push the pretrained generation model away from its training distribution, leading to degraded generation quality and poor compatibility with other extension models. Therefore, there is a pressing need to propose a new end-to-end framework that enables consistent character generation.
@@
Here we are particularly interested in consistent image generation for human. Since text-to-image models are pretrained on large-scale image-text data, which contains massive text prompts with celeb names, the models can generate identity-consistent images using celeb names. These names are ideal examples for this task. Previous work (wang2024stableidentity, ) has revealed that the word embeddings of celeb names constitute a human-centric prior space with editability, so we decide to conduct new character sampling in this space.
@@
In this work, we propose CharacterFactory, a framework for new character creation which mainly consists of an Identity-Embedding GAN (IDE-GAN) and a context-consistent loss. Specifically, a GAN model composed of MLPs is used to map from a latent space to the celeb embedding space following the adversarial learning manner, with word embeddings of celeb names as real data and generated ones as fake. Furthermore, to enable the generated embeddings to work like the native word embeddings of CLIP (radford2021learning, ), we constrain these embeddings to exhibit consistency when combined with diverse contexts. Following this paradigm, the generated embeddings could be naturally inserted into CLIP text encoder, hence could be seamlessly integrated with the image/video/3D diffusion models. In addition, since IDE-GAN is composed of only MLPs as trainable parameters and accesses only the pretrained CLIP during training, it takes only 10 minutes to train and then infinite new identity embeddings could be sampled to produce identity-consistent images for new characters during inference.
@@ 2.1. Text-to-Image Diffusion Models
Recent advances in diffusion models (ho2020denoising, ; song2020denoising, ) have shown unprecedented capabilities for text-to-image generation (rombach2022high, ; nichol2021glide, ; ramesh2022hierarchical, ), and new possibilities are still emerging (wang2022hs, ; chen2024shoemodel, ). The amazing generation performance is derived from the high-quality large-scale image-text pairs (schuhmann2022laion, ; schuhmann2021laion, ), flourishing foundational models (chen2023pixart, ; peebles2023scalable, ), and stronger controllability design (zhang2023adding, ; zhang2024transparent, ). Their fundamental principles are based on Denoising Diffusion Probabilistic Models (DDPMs) (ho2020denoising, ), which include a forward noising process and a reverse denoising process. The forward process adds Gaussian noise progressively to an input image, and the reverse process is modeled with a UNet trained for predicting noise. Supervised by the denoising loss, a random noise can be denoised to a realistic image by iterating the reverse diffusion process. However, due to the stochastic nature of this generation process, existing text-to-image diffusion models are not able to directly implement consistent character generation.
@@ 2.2. Consistent Character Generation
Existing works on consistent character generation mainly focus on personalization for the target subject (gal2022image, ; ruiz2023dreambooth, ). Textual Inversion (gal2022image, ) represents the target subject as a new word embedding via optimization while freezing the diffusion model. DreamBooth (ruiz2023dreambooth, ) finetunes all weights of the diffusion model to fit only the target subject. IP-Adapter (ye2023ip, ) designs a decoupled cross-attention mechanism for text features and image features. Celeb-Basis (yuan2023inserting, ) and StableIdentity (wang2024stableidentity, ) use prior information from celeb names to make optimization easier and improve editability. PhotoMaker trains MLPs and the LoRA residuals of the attention layers to inject identity information (li2023photomaker, ). But these methods attempt to produce identity-consistent images based on the reference images, instead of creating a new character. In addition, The Chosen One (avrahami2023chosen, ) clusters the generated images to obtain similar outputs for learning a customized model on a highly similar cluster by iterative optimization with personalized LoRA weights and word embeddings, which is a time-consuming process. ConsiStory (tewel2024training, ) introduces a shared attention block mechanism and correspondence-based feature injection between a batch of images, but relying only on patch features lacks semantic understanding for the subject and makes the inference process complicated. Despite creating new characters, they still suffer from complicated pipelines and poor editability.
@@ 2.3. Integrating Diffusion Models and GANs
Generative Adversarial Net (GAN) (goodfellow2014generative, ; karras2019style, ) models the mapping between data distributions by adversarially training a generator and a discriminator. Although GAN-based methods have been outperformed by powerful diffusion models for image generation, they perform well on small-scale datasets (dhariwal2021diffusion, ) benefiting from the flexibility of GANs. Some methods focus on combining them to improve the optimization objective for diffusion models with GANs (xu2023ufogen, ; xiao2021tackling, ; wang2022diffusion, ). In this work, we for the first time construct a GAN model in CLIP embedding space to sample consistent identity for diffusion models.
@@ 3.1. Preliminary
In this work, we employ the pretrained Stable Diffusion (rombach2022high, ) (denoted as SD) as the base text-to-image model. SD consists of three components: a CLIP text encoder $e_{text}$ (radford2021learning, ), a Variational Autoencoder (VAE) ($\mathcal{E}$, $\mathcal{D}$(esser2021taming, ) and a denoising U-Net $\epsilon_{\theta}$. With the text conditioning, $\epsilon_{\theta}$ can denoise sampled Gaussian noises to realistic images conforming to the given text prompts $p$. In particular, the tokenizer of $e_{text}$ sequentially divides and encodes $p$ into $l$ integer tokens. Subsequently, by looking up the tokenizer’s dictionary, the embedding layer of $e_{text}$ retrieves a group of corresponding word embeddings $g=[v_{1},...,v_{l}],v_{i}\in\mathbb{R}^{d}$. Then, the text transformer $\tau_{text}$ of $e_{text}$ further represents $g$ to contextual embeddings $\bar{g}=[\bar{v}_{1},...,\bar{v}_{l}],\bar{v}_{i}\in\mathbb{R}^{d}$ with the cross-attention mechanism. And $\epsilon_{\theta}$ renders the content conveyed in text prompts by cross attention between $\bar{g}$ and diffusion features.
@@ 3.2. IDE-GAN
Specifically, we employ 326 celeb names (wang2024stableidentity, ) which consist only of first name and last name, and encode them into the corresponding word embeddings $C\in\mathbb{R}^{326\times 2\times d}$ for training. In addition, we observe that adding a small noise to the celeb embeddings can still generate images with corresponding identity. Therefore, we empirically introduce random noise $\eta\sim\mathcal{N}(0,\mathbf{I})$ scaled by $35e-$ 3 as a data augmentation. As shown in Figure 2(a), given a latent code $z\in\mathcal{N}(0,\mathbf{I})$, the generator $G$ is trained to produce embeddings $[v_{1}^{*},v_{2}^{*}]$ that cannot be distinguished from “real” (i.e., celeb embeddings) by an adversarially trained discriminator $D$. To alleviate the training difficulty of $G$, we use AdaIN to help the MLPs’ output embeddings $[v^{\prime}_{1},v^{\prime}_{2}]$ land more naturally into the celeb embedding space (wang2024stableidentity, ):
@@ 3.3. Context-Consistent Loss
Specifically, we sample 1,000 text prompts with ChatGPT (ouyang2022training, ) for various contexts (covering expressions, decorations, actions, attributes, and backgrounds), like “Under the tree, $s_{1}^{*}$ $s_{2}^{*}$ has a picnic”, and demand that the position of “$s_{1}^{*}$ $s_{2}^{*}$” in the context should be as diverse as possible. During training, we sample $N$ prompts from the collected prompt set, and use the tokenizer and embedding layer to encode them into $N$ groups of word embeddings. The generated embeddings $[v_{1}^{*},v_{2}^{*}]$ are inserted at the position of “$s_{1}^{*}$ $s_{2}^{*}$”. Then, the text transformer $\tau_{text}$ further represents them to $N$ groups of contextual embeddings, where we expect to minimize the average pairwise distances among the $\{[\bar{v}_{1}^{*},\bar{v}_{2}^{*}]_{i}\}_{i=1}^{N}$:
@@ 3.4. Full Objective
  • GAN for word embedding. We introduce GAN in the CLIP embedding space for the first time and leverage the subsequent network to design the context-consistent loss which can perceive the generated pseudo identity embeddings in diverse contexts. This design is similar to the thought of previous works for image generation (isola2017image, ; zhang2018unreasonable, ; cao2023multi, ), which have demonstrated that mixing the GAN objective and a more traditional loss such as L2 distance is beneficial.
@@
Figure 4. Qualitative comparisons with five baselines in diverse text prompts (covering actions, attributes, decorations and a style). To avoid considering identity preservation with the input images, we do not show the input and use different identities for each baseline. Baselines either perform poorly in identity consistency or editability, and the proposed CharacterFactory is able to maintain consistent identities while conforming to the text prompts.
@@ 4.1. Experimental Setting
Implementation Details. We employ Stable Diffusion v2.1-base as our base model. The number of layers in the MLPs for the generator $G$ and the discriminator $D$ are 2 and 3 respectively. The dimension of $z$ is set to 64 empirically. The batch size and learning rate are set to 1 and $55e-$ 5. We employ an Adam optimizer (kingma2014adam, ) with the momentum parameters $\beta_{1}=0.5$ and $\beta_{2}=0.999$ to optimize our IDE-GAN. The trade-off parameters $\lambda_{1}$ and $\lambda_{2}$ are both 1 as default. CharacterFactory is trained with only 10 minutes for 10,000 steps on a single NVIDIA A100. The classifier-free guidance (ho2022classifier, ) scale is 8.5 for inference as default. More implementation details can be found in the supplementary material.
Baselines. Since the most related methods, The Chosen One (avrahami2023chosen, ) and ConsiStory (tewel2024training, ) which are also designed for consistent text-to-image generation, have not released their codes yet, we compare these methods with the content provided in their papers. In addition, we also compare with the subject-driven generation methods (related techniques) which include the optimization-based methods: Textual Inversion (gal2022image, ), DreamBooth (ruiz2023dreambooth, ), Celeb-Basis (yuan2023inserting, ), and the encoder-based methods: IP-Adapter (ye2023ip, ), PhotoMaker (li2023photomaker, ). We prioritize to use the official models released by these methods. We use the Stable Diffusion 2.1 versions of Textual Inversion and DreamBooth for fair comparison.
Evaluation. The input of our method comes from random noise, so this work does not compare subject preservation. But subject-driven generation methods need input images, so we randomly select 70 non-celeb images from FFHQ (karras2019style, ) for these methods. These input images are used only for subject information injection and not shown in qualitative comparisons, nor involved in the calculation of quantitative comparisons. To conduct a comprehensive evaluation, we use 40 text prompts that cover decorations, actions, expressions, attributes and backgrounds (li2023photomaker, ). Overall, we use 70 identities and 40 text prompts to generate 2,800 images for each competing method.
Table 1. Quantitative comparisons with five baselines. $\uparrow$ indicates higher is better, and $\downarrow$ indicates that lower is better. The best results are shown in bold. The proposed CharacterFactory obtains superior identity consistency, editability, trusted face diversity and image quality than the subject-driven generation methods.
@@
Textual Inversion (gal2022image, ) 0.654 0.379 0.282 0.370 0.108 28.64
DreamBooth (ruiz2023dreambooth, ) 0.695 0.462 0.289 0.367 0.059 48.55
IP-Adapter (ye2023ip, ) 0.872 0.468 0.236 0.256 0.111 78.95
Celeb-Basis (yuan2023inserting, ) 0.673 0.400 0.278 0.379 0.134 33.72
PhotoMaker (li2023photomaker, ) 0.717 0.465 0.291 0.304 0.126 41.38
CharacterFactory 0.764 0.498 0.332 0.333 0.140 22.58
@@
Figure 5. Qualitative comparisons with the generation results in the papers of two most related methods The Chosen One (avrahami2023chosen, ) and ConsiStory (tewel2024training, ). CharacterFactory achieves comparable performance with the same prompts (Zoom in for the best view).
@@
Metrics: We calculate the CLIP visual similarity (CLIP-I) between the generated results of “a photo of $s_{1}^{*}$ $s_{2}^{*}$” and other text prompts to evaluate Subject Consistency. And we calculate face similarity (deng2019arcface, ) and perceptual similarity (i.e., LPIPS) (zhang2018perceptual, ) between the detected face regions with the same settings to measure the Identity Consistency and Face Diversity (li2023photomaker, ; wu2023singleinsert, ). But inconsistent faces might obtain high face diversity, leading to unreliable results. Therefore, we also introduce the Trusted Face Diversity (wang2024stableidentity, ) which is calculated by the product of cosine distances from face similarity and face diversity between each pair of images, to evaluate whether the generated faces from the same identity are both consistent and diverse. We calculate the text-image similarity (CLIP-T) to measure the Editablity. In addition, we randomly sample 70 celeb names to generate images with the introduced 40 text prompts as pseudo ground truths, and calculate Fréchet Inception Distance (FID) (lucic2017gans, ) between the generated images by competing methods and pseudo ground truths to measure the Image Quality.
4.2. Comparison with Subject-Driven Methods
@@
Qualitative Comparison. As mentioned in Section 4.1, we randomly select 70 images from FFHQ (karras2019style, ) to inject identity information for subject-driven baselines, and mainly compare identity consistency and editability, as shown in Figure 4. PhotoMaker and Celeb-Basis are human-centric methods. The former pretrains a face encoder and LoRA residuals on large-scale datasets, but still suffers from identity deviation for editing (Column 6, 8). The latter optimizes word embeddings to represent the target identity, but fails to edit with “white hair” and produces a fatter face (Column 3, 7). IP-Adapter (ye2023ip, ) learns text-image decoupled cross attention but still struggles to present “white hair” and accurate “doctoral cap”, and cannot be stylized with text prompt (Column 3, 6, 8). DreamBooth (ruiz2023dreambooth, ) finetunes the whole model to adapt to the input image and tends to generate images similar to the input image. It lacks layout diversity and fails to produce “wearing a blue hoodie”. Due to the stochasticity of Textual Inversion (gal2022image, )’s optimization process, its identity consistency and image quality are relatively weak. Overall, baselines show decent performance for identity consistency and editability, and both rely on the input image. In contrast, the proposed CharacterFactory can sample pseudo identities end-to-end and generate identity-consistent prompt-aligned results.
@@
The Chosen One (avrahami2023chosen, ) 1,200 LoRAs + two word embeddings
@@
Quantitative Comparison. In addition, we also provide the quantitative comparison with five baselines in Table 1. Since IP-Adapter tends to generate frontal faces, it obtains better subject consistency (CLIP-I) but weak editability (CLIP-T). CLIP-I mainly measures high-level semantic alignment and lacks the assessment for identity, so we further introduce the identity consistency for evaluation. Our method achieves the best identity consistency, editability and second-place subject consistency. In particular, the proposed context-consistent loss incentivizes pseudo identities to exhibit consistency in various contexts. On the other hand, our effective adversarial learning enables pseudo identity embeddings to work in Stable Diffusion as naturally as celeb embeddings, and thus outperforms PhotoMaker (the second place) by 0.041 on editability. Textual Inversion and DreamBooth obtain good face diversity but weak trusted diversity. This is because face diversity measures whether the generated faces from the same identity are diverse in different contexts, but inconsistent identities can also be incorrectly recognized as “diverse”. Therefore, trusted face diversity is introduced to evaluate whether the results are both consistent and diverse. So the Celeb-Basis obtains the best face diversity. but inferior to the proposed CharacterFactory 0.006 on trusted face diversity. For the image quality (FID), Textual Inversion, Celeb-Basis and ours without finetuning Stable Diffusion can maintain more native generation performance, and obtain better image quality. Since our generated pseudo identity embeddings are closer to the distribution of celeb embeddings, producing more natural generation results and the best image quality ($+6.06$ than the second place).

### v2

@@ 1. Introduction
In the evolving realm of text-to-image generation, diffusion models have emerged as indispensable tools for content creation (Rombach et al., 2022; Chen et al., 2023b; Zhang et al., 2023c). However, the inherent stochastic nature of the generation models leads to the inability to generate consistent subjects in different contexts directly, as shown in Figure 1. Such consistency can derive many applications: illustrating books and stories, creating brand ambassador, movie making, developing presentations, art design, identity-consistent data construction and more.
@@
Subject-driven methods work by either representing a user-specific image as a new word (Li et al., 2023; Wang et al., 2024b; Chen et al., 2023a) or learning image feature injection (Wang et al., 2024a; Ye et al., 2023; Wei et al., 2023) for consistent image generation. Their training paradigms typically include per-subject optimization and encoder pretraining on large-scale datasets. The former usually requires lengthy optimization for each subject and tends to overfit the appearance in the input image (Gal et al., 2022; Ruiz et al., 2023). The latter consumes significant computational costs and struggles in stably capturing the identity and its details (Li et al., 2023; Wang et al., 2024a). However, these methods attempt to produce images with the same identity as the reference images, instead of creating a new character in various contexts. A feasible way is that a text-to-image model is used in advance to create a new character’s image and then subject-driven methods are adopted to produce images with consistent identity. Such a two-stage workflow could push the pretrained generation model away from its training distribution, leading to degraded generation quality and poor compatibility with other extension models. Therefore, there is a pressing need to propose a new end-to-end framework that enables consistent character generation.
@@
Here we are particularly interested in consistent image generation for human. Since text-to-image models are pretrained on large-scale image-text data, which contains massive text prompts with celeb names, the models can generate identity-consistent images using celeb names. These names are ideal examples for this task. Previous work (Wang et al., 2024b) has revealed that the word embeddings of celeb names constitute a human-centric prior space with editability, so we decide to conduct new character sampling in this space.
@@
In this work, we propose CharacterFactory, a framework for new character creation which mainly consists of an Identity-Embedding GAN (IDE-GAN) and a context-consistent loss. Specifically, a GAN model composed of MLPs is used to map from a latent space to the celeb embedding space following the adversarial learning manner, with word embeddings of celeb names as real data and generated ones as fake. Furthermore, to enable the generated embeddings to work like the native word embeddings of CLIP (Radford et al., 2021), we constrain these embeddings to exhibit consistency when combined with diverse contexts. Following this paradigm, the generated embeddings could be naturally inserted into CLIP text encoder, hence could be seamlessly integrated with the image/video/3D diffusion models. In addition, since IDE-GAN is composed of only MLPs as trainable parameters and accesses only the pretrained CLIP during training, it takes only 10 minutes to train and then infinite new identity embeddings could be sampled to produce identity-consistent images for new characters during inference.
@@ 2.1. Text-to-Image Diffusion Models
Recent advances in diffusion models (Ho et al., 2020; Song et al., 2020) have shown unprecedented capabilities for text-to-image generation (Rombach et al., 2022; Nichol et al., 2021; Ramesh et al., 2022), and new possibilities are still emerging (Wang et al., 2022a; Chen et al., 2024). The amazing generation performance is derived from the high-quality large-scale image-text pairs (Schuhmann et al., 2022, 2021), flourishing foundational models (Chen et al., 2023b; Peebles and Xie, 2023), and stronger controllability design (Zhang et al., 2023b; Zhang and Agrawala, 2024). Their fundamental principles are based on Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020), which include a forward noising process and a reverse denoising process. The forward process adds Gaussian noise progressively to an input image, and the reverse process is modeled with a UNet trained for predicting noise. Supervised by the denoising loss, a random noise can be denoised to a realistic image by iterating the reverse diffusion process. However, due to the stochastic nature of this generation process, existing text-to-image diffusion models are not able to directly implement consistent character generation.
@@ 2.2. Consistent Character Generation
Existing works on consistent character generation mainly focus on personalization for the target subject (Gal et al., 2022; Ruiz et al., 2023). Textual Inversion (Gal et al., 2022) represents the target subject as a new word embedding via optimization while freezing the diffusion model. DreamBooth (Ruiz et al., 2023) finetunes all weights of the diffusion model to fit only the target subject. IP-Adapter (Ye et al., 2023) designs a decoupled cross-attention mechanism for text features and image features. Celeb-Basis (Yuan et al., 2023) and StableIdentity (Wang et al., 2024b) use prior information from celeb names to make optimization easier and improve editability. PhotoMaker trains MLPs and the LoRA residuals of the attention layers to inject identity information (Li et al., 2023). But these methods attempt to produce identity-consistent images based on the reference images, instead of creating a new character. In addition, The Chosen One (Avrahami et al., 2023) clusters the generated images to obtain similar outputs for learning a customized model on a highly similar cluster by iterative optimization with personalized LoRA weights and word embeddings, which is a time-consuming process. ConsiStory (Tewel et al., 2024) introduces a shared attention block mechanism and correspondence-based feature injection between a batch of images, but relying only on patch features lacks semantic understanding for the subject and makes the inference process complicated. Despite creating new characters, they still suffer from complicated pipelines and poor editability.
@@ 2.3. Integrating Diffusion Models and GANs
Generative Adversarial Net (GAN) (Goodfellow et al., 2014; Karras et al., 2019) models the mapping between data distributions by adversarially training a generator and a discriminator. Although GAN-based methods have been outperformed by powerful diffusion models for image generation, they perform well on small-scale datasets (Dhariwal and Nichol, 2021) benefiting from the flexibility of GANs. Some methods focus on combining them to improve the optimization objective for diffusion models with GANs (Xu et al., 2023; Xiao et al., 2021; Wang et al., 2022b). In this work, we for the first time construct a GAN model in CLIP embedding space to sample consistent identity for diffusion models.
@@ 3.1. Preliminary
In this work, we employ the pretrained Stable Diffusion (Rombach et al., 2022) (denoted as SD) as the base text-to-image model. SD consists of three components: a CLIP text encoder $e_{text}$ (Radford et al., 2021), a Variational Autoencoder (VAE) ($\mathcal{E}$, $\mathcal{D}$(Esser et al., 2021) and a denoising U-Net $\epsilon_{\theta}$. With the text conditioning, $\epsilon_{\theta}$ can denoise sampled Gaussian noises to realistic images conforming to the given text prompts $p$. In particular, the tokenizer of $e_{text}$ sequentially divides and encodes $p$ into $l$ integer tokens. Subsequently, by looking up the tokenizer’s dictionary, the embedding layer of $e_{text}$ retrieves a group of corresponding word embeddings $g=[v_{1},...,v_{l}],v_{i}\in\mathbb{R}^{d}$. Then, the text transformer $\tau_{text}$ of $e_{text}$ further represents $g$ to contextual embeddings $\bar{g}=[\bar{v}_{1},...,\bar{v}_{l}],\bar{v}_{i}\in\mathbb{R}^{d}$ with the cross-attention mechanism. And $\epsilon_{\theta}$ renders the content conveyed in text prompts by cross attention between $\bar{g}$ and diffusion features.
@@ 3.2. IDE-GAN
Specifically, we employ 326 celeb names (Wang et al., 2024b) which consist only of first name and last name, and encode them into the corresponding word embeddings $C\in\mathbb{R}^{326\times 2\times d}$ for training. In addition, we observe that adding a small noise to the celeb embeddings can still generate images with corresponding identity. Therefore, we empirically introduce random noise $\eta\sim\mathcal{N}(0,\mathbf{I})$ scaled by $35e-$ 3 as a data augmentation. As shown in Figure 2(a), given a latent code $z\in\mathcal{N}(0,\mathbf{I})$, the generator $G$ is trained to produce embeddings $[v_{1}^{*},v_{2}^{*}]$ that cannot be distinguished from “real” (i.e., celeb embeddings) by an adversarially trained discriminator $D$. To alleviate the training difficulty of $G$, we use AdaIN to help the MLPs’ output embeddings $[v^{\prime}_{1},v^{\prime}_{2}]$ land more naturally into the celeb embedding space (Wang et al., 2024b):
@@ 3.3. Context-Consistent Loss
Specifically, we sample 1,000 text prompts with ChatGPT (Ouyang et al., 2022) for various contexts (covering expressions, decorations, actions, attributes, and backgrounds), like “Under the tree, $s_{1}^{*}$ $s_{2}^{*}$ has a picnic”, and demand that the position of “$s_{1}^{*}$ $s_{2}^{*}$” in the context should be as diverse as possible. During training, we sample $N$ prompts from the collected prompt set, and use the tokenizer and embedding layer to encode them into $N$ groups of word embeddings. The generated embeddings $[v_{1}^{*},v_{2}^{*}]$ are inserted at the position of “$s_{1}^{*}$ $s_{2}^{*}$”. Then, the text transformer $\tau_{text}$ further represents them to $N$ groups of contextual embeddings, where we expect to minimize the average pairwise distances among the $\{[\bar{v}_{1}^{*},\bar{v}_{2}^{*}]_{i}\}_{i=1}^{N}$:
@@ 3.4. Full Objective
  • GAN for word embedding. We introduce GAN in the CLIP embedding space for the first time and leverage the subsequent network to design the context-consistent loss which can perceive the generated pseudo identity embeddings in diverse contexts. This design is similar to the thought of previous works for image generation (Isola et al., 2017; Zhang et al., 2018a; Cao et al., 2023), which have demonstrated that mixing the GAN objective and a more traditional loss such as L2 distance is beneficial.
@@
Figure 4. Qualitative comparisons with two-stage workflows using five baselines (denoted with $\dagger$) for creating consistent characters. The upper left corner of the two-stage baselines is the generated image by Stable Diffusion as the input of the second stage. Two-stage workflows struggle to maintain the identity of the generated image and degrade the image quality. In comparison, the proposed CharacterFactory can generate high-quality identity-consistent character images with diverse layouts while conforming to the given text prompts (Zoom in for the best view).
@@ 4.1. Experimental Setting
Implementation Details. We employ Stable Diffusion v2.1-base as our base model. The number of layers in the MLPs for the generator $G$ and the discriminator $D$ are 2 and 3 respectively. The dimension of $z$ is set to 64 empirically. The batch size and learning rate are set to 1 and $55e-$ 5. We employ an Adam optimizer (Kingma and Ba, 2014) with the momentum parameters $\beta_{1}=0.5$ and $\beta_{2}=0.999$ to optimize our IDE-GAN. The trade-off parameters $\lambda_{1}$ and $\lambda_{2}$ are both 1 as default. CharacterFactory is trained with only 10 minutes for 10,000 steps on a single NVIDIA A100. The classifier-free guidance (Ho and Salimans, 2022) scale is 8.5 for inference as default. More implementation details can be found in the supplementary material.
Table 1. Quantitative comparisons with two-stage workflows using five baselines (denoted with $\dagger$). $\uparrow$ indicates higher is better, and $\downarrow$ indicates that lower is better. The best results are shown in bold. We define the speed as the time it takes to create a new consistent character on a single NVIDIA A100 GPU. Obviously, CharacterFactory obtains superior performance on identity consistency, editability, trusted face diversity, image quality and speed, which are consistent with the qualitative comparisons.
@@
Textual Inversion (Gal et al., 2022) 0.647 0.295 0.274 0.392 0.078 47.94 3200
DreamBooth (Ruiz et al., 2023) 0.681 0.443 0.287 0.339 0.073 62.66 1500
IP-Adapter (Ye et al., 2023) 0.853 0.447 0.227 0.192 0.096 95.25 7
Celeb-Basis (Yuan et al., 2023) 0.667 0.369 0.273 0.378 0.101 56.43 480
PhotoMaker (Li et al., 2023) 0.694 0.451 0.301 0.331 0.138 53.37 10
CharacterFactory 0.764 0.498 0.332 0.333 0.140 22.58 3
@@
Figure 5. Qualitative comparisons with the generation results in the papers of two most related methods The Chosen One (Avrahami et al., 2023) and ConsiStory (Tewel et al., 2024). CharacterFactory achieves comparable performance with the same prompts (Zoom in for the best view).
Baselines. Since the most related methods, The Chosen One (Avrahami et al., 2023) and ConsiStory (Tewel et al., 2024) which are also designed for consistent text-to-image generation, have not released their codes yet, we compare these methods with the content provided in their papers. In addition, as we introduced in Section 1, the two-stage workflows with subject-driven methods can also create new characters. Therefore, we first use a prompt “a photo of a person, facing to the camera” to drive Stable Diffusion to generate images of new characters as the input of the second stage, and then use these subject-driven methods to produce character images with diverse prompts for comparison. These input images are used for subject information injection and not involved in the calculation of quantitative comparisons. These methods include the optimization-based methods: Textual Inversion (Gal et al., 2022), DreamBooth (Ruiz et al., 2023), Celeb-Basis (Yuan et al., 2023), and the encoder-based methods: IP-Adapter (Ye et al., 2023), PhotoMaker (Li et al., 2023). We prioritize to use the official models released by these methods. We use the Stable Diffusion 2.1 versions of Textual Inversion and DreamBooth for fair comparison.
Evaluation. The input of our method comes from random noise, so this work does not compare subject preservation for quantitative comparison. To conduct a comprehensive evaluation, we use 40 text prompts that cover decorations, actions, expressions, attributes and backgrounds (Li et al., 2023). Overall, we use 70 identities and 40 text prompts to generate 2,800 images for each competing method.
@@
Metrics: We calculate the CLIP visual similarity (CLIP-I) between the generated results of “a photo of $s_{1}^{*}$ $s_{2}^{*}$” and other text prompts to evaluate Subject Consistency. And we calculate face similarity (Deng et al., 2019) and perceptual similarity (i.e., LPIPS) (Zhang et al., 2018b) between the detected face regions with the same settings to measure the Identity Consistency and Face Diversity (Li et al., 2023; Wu et al., 2023). But inconsistent faces might obtain high face diversity, leading to unreliable results. Therefore, we also introduce the Trusted Face Diversity (Wang et al., 2024b) which is calculated by the product of cosine distances from face similarity and face diversity between each pair of images, to evaluate whether the generated faces from the same identity are both consistent and diverse. We calculate the text-image similarity (CLIP-T) to measure the Editablity. In addition, we randomly sample 70 celeb names to generate images with the introduced 40 text prompts as pseudo ground truths, and calculate Fréchet Inception Distance (FID) (Lucic et al., 2017) between the generated images by competing methods and pseudo ground truths to measure the Image Quality.
4.2. Comparison with Two-Stage Workflows.
@@
Qualitative Comparison. As mentioned in Section 4.1, we randomly generate 70 character images in front view to inject identity information for two-stage workflows using subject-driven baselines (denoted with $\dagger$), as shown in Figure 4. PhotoMaker (Li et al., 2023) and Celeb-Basis (Yuan et al., 2023) are human-centric methods. The former pretrains a face encoder and LoRA residuals on large-scale datasets. The latter optimizes word embeddings to represent the target identity. But they all suffer from degraded image quality under this setting. IP-Adapter (Ye et al., 2023) learns text-image decoupled cross attention, but fails to present “Christmas hat” and “spacesuit”. DreamBooth (Ruiz et al., 2023) finetunes the whole model to adapt to the input image and tends to generate images similar to the input image. It lacks generation diversity and fails to produce the “Christmas hat”. Due to the stochasticity of Textual Inversion (Gal et al., 2022)’s optimization process, its identity consistency and image quality are relatively weak. Overall, two-stage workflows show decent performance for identity consistency, editability, and image quality, and they all rely on the input images and struggle to preserve the input identity. In contrast, the proposed CharacterFactory can sample pseudo identities end-to-end and generate identity-consistent prompt-aligned results with high quality.
@@
The Chosen One (Avrahami et al., 2023) 1,200 LoRAs + two word embeddings
@@
Quantitative Comparison. In addition, we also provide the quantitative comparison with five baselines in Table 1. Since IP-Adapter tends to generate frontal faces, it obtains better subject consistency (CLIP-I) but weak editability (CLIP-T). CLIP-I mainly measures high-level semantic alignment and lacks the assessment for identity, so we further introduce the identity consistency for evaluation. Our method achieves the best identity consistency, editability and second-place subject consistency. In particular, the proposed context-consistent loss incentivizes pseudo identities to exhibit consistency in various contexts. On the other hand, our effective adversarial learning enables pseudo identity embeddings to work in Stable Diffusion as naturally as celeb embeddings, and thus outperforms PhotoMaker (the second place) by 0.031 on editability. Textual Inversion and Celeb-Basis obtain good face diversity but weak trusted diversity. This is because face diversity measures whether the generated faces from the same identity are diverse in different contexts, but inconsistent identities can also be incorrectly recognized as “diverse”. Therefore, trusted face diversity is introduced to evaluate whether the results are both consistent and diverse. So Textual Inversion obtains the best face diversity, but is inferior to CharacterFactory 0.062 on trusted face diversity. For image quality (FID), the two-stage workflows directly lead to an unacceptable degradation of competing methods on image quality quantitatively. On the other hand, two-stage workflows consume more time for creating identity-consistent characters. In comparison, our end-to-end framework implements more natural generation results, the best image quality and faster inference workflow.